# GEIC: Universal and Multilingual Named Entity Recognition with Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have supplanted traditional methods in numerous natural language processing tasks. Nonetheless, in Named Entity Recognition (NER), existing LLM-based methods underperform compared to baselines and require significantly more computational resources, limiting their application. In this paper, we introduce the task of generation-based extraction and in-context classification (GEIC), designed to leverage LLMs’ prior knowledge and self-attention mechanisms for NER tasks. We then propose CascadeNER, a universal and multilingual GEIC framework for few-shot and zero-shot NER. CascadeNER employs model cascading to utilize two small-parameter LLMs to extract and classify independently, reducing resource consumption while enhancing accuracy. We also introduce AnythingNER, the first NER dataset specifically designed for LLMs, including 8 languages, 155 entity types and a novel dynamic categorization system. Experiments show that CascadeNER achieves state-of-the-art performance on low-resource and fine-grained scenarios, including CrossNER and FewNERD. Our work is openly accessible.
GEIC: Universal and Multilingual Named Entity Recognition
with Large Language Models
Hanjun Luothanks: Zhejiang University, hanjun.21@intl.zju.edu.cn 11footnotemark: 1, Yingbin Jinthanks: Zhejiang University, 3210104412@zju.edu.cn 11footnotemark: 1, Xuecheng Liuthanks: Zhejiang University, xuecheng.21@intl.zju.edu.cn 11footnotemark: 1, Tong Shangthanks: Southeast University, 213211756@seu.edu.cn 22footnotemark: 2, Ruizhe Chenthanks: Zhejiang University, ruizhec.21@intl.zju.edu.cn 11footnotemark: 1, Zuozhu Liuthanks: Zhejiang University, zuozhuliu@intl.zju.edu.cn, Corresponding author 11footnotemark: 1
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has revolutionized natural language processing (NLP) (Zhao et al., 2023; Naveed et al., 2023). Recent LLMs, such as GPT-4 (Achiam et al., 2023) and Claude-3 Anthropic (2023), with their powerful generalization and attention mechanisms Vaswani et al. (2017), excel not only in text generation tasks like Text Summarization Wilson et al. (2024); Jin et al. (2024), Question Answering Tan et al. (2023), and Machine Translation Vilar et al. (2022); Moslem et al. (2023), but also in discriminative tasks, including Named Entity Extraction Sancheti et al. (2024) and Text Classification Gasparetto et al. (2022).
Nevertheless, unlike other discriminative tasks, existing LLM-based approaches Shao et al. (2023); Keloth et al. (2024); Li and Zhang (2023) struggle to replace supervised methods in Named Entity Recognition (NER). NER is typically framed as a sequence labeling task, where the model assigns entity labels to each token in a sentence. This structure emphasizes sequential dependencies, which contrasts with LLMs’ focus on global attention mechanisms. Although existing methods leverage generalization and prior knowledge of LLMs to outperform supervised baselines in low-resource scenarios, they often fall short significantly in standard scenarios or complex entity classifications. Moreover, their reliance on commercial models, typically GPT, introduces significant challenges related to cost, data security, and transparency, limiting their broader applicability.
To address these issues, we introduce Generation-based Extraction and In-context Classification (GEIC), a novel task designed to harness LLMs’ attention mechanisms and prior knowledge for NER. As illustrated in Figure 1, unlike traditional sequence labeling methods, GEIC redefines NER as two sequential in-context text generation sub-tasks, extraction and classification, which decreases the task complexity and enhances the capability to capture in-context dependencies. To achieve GEIC, we propose CascadeNER, a universal and multilingual GEIC framework. CascadeNER employs model cascading Varshney and Baral (2022), assigning the each sub-tasks to corresponding fine-tuned small-parameter LLMs, referred to as Small Language Models (SLMs). This method reduces computational resource consumption while maintaining high performance, enabling efficient NER with SLMs. Leveraging the prior knowledge, CascadeNER can complete NER tasks in multiple languages in both few-shot and zero-shot scenarios. Extensive experiments demonstrate that CascadeNER achieves state-of-the-art (SOTA) performance on low-resource and fine-grained datasets. Considering the lack of datasets designed for LLM-based NER, we collect AnythingNER for CascadeNER, a multilingual and fine-grained NER dataset, featuring 8 major languages, 8 coarse-grained types, 31 medium-grained types, 155 fine-grained types, and a novel dynamic categorization system. As the first dataset optimized for LLMs’ generalization and few-shot/zero-shot learning in NER, AnythingNER provides an unprecedented level of detail and coverage. Moreover, this work offers the first comprehensive comparison and analysis of existing LLM-based NER methods, with a emphasis on multilingual and fine-grained scenarios.
Figure 1: Red arrows indicate that the data is used for training. Traditional methods sequentially label each token according to predefined categories from the train set. GEIC does not differentiate between entity types when extracting but introduces a classification sub-task, thereby improving its generalization. The categories can be changed without re-training, enabling zero-shot reasoning.
@@
  • We introduce generation-based extraction and in-context classification (GEIC), a novel task to replace traditional NER tasks in LLMs, emphasizing global attention and generalization
  • We propose CascadeNER, the first universal and multilingual NER framework with SLMs, which supports both few-shot and zero-shot scenarios and achieves SOTA performance on low-resource and fine-grained datasets
  • We build AnythingNER, the first multilingual and fine-grained datasets designed for NER with LLMs, especially GEIC, with a novel dynamic categorization system
  • We conduct comprehensive comparison of existing LLM-based NER methods, providing future directions of LLM-based NER
@@ 2 Related Works
2.1 Named Entity Recognition
@@
Named Entity Recognition (NER) is the task of identifying named entities in text and classifying it into predefined categories. Supervised methods, such as BiLSTM Yu et al. (2020) and BERT-MRC Li et al. (2019a), currently dominate this task. They generally rely on large amounts of train data to achieve strong performance, which limits their application in low-resource scenarios. With the advancement of LLMs, leveraging their few-shot and zero-shot learning capabilities to address this. Xie et al. (2023) explore the application of GPT-3 Brown (2020) for NER, but its performance is significantly lower than supervised methods’. GPT-NER Wang et al. (2023) improve GPT-3’s NER capabilities by re-framing the task as single-entity labeling through prompt engineering. While GPT-NER enhanced performance, it still lags behind supervised methods. PromptNER Ashok and Lipton (2023) achieves better results with GPT-4, yet failed to deliver a robust zero-shot solution. Furthermore, several studies investigate LLM-based NER in domain-specific tasks Li and Zhang (2023); Shao et al. (2023); Keloth et al. (2024). These studies, however, all rely on commercial models, increasing costs and limiting flexibility. Existing results highlight both the potential of LLMs in NER and the pressing need for more adaptable and cost-effective solutions.
2.2 Small Language Models
While commercial LLMs undergoes rapid development, small-parameter and open-source LLMs, known as small language models (SLMs), also attracts significant attention Touvron et al. (2023); Hu et al. (2024a); Zhang et al. (2024); GLM et al. (2024); Bai et al. (2023). These models employ techniques such as quantization Han et al. (2015) and distillation Hinton (2015) to reduce their parameters to between 1.5B and 7B,offering advantages like easier fine-tuning and lower computational resource consumption, albeit at the cost of some performance degradation in complex tasks. This trade-off have led to the widely use of SLMs like LLaMA-2-7B in both academia and industry Clemmer et al. (2024); Gade et al. (2023); Luo et al. (2024). However, NER, as a relatively complex task, places high demands on model performance. Existing LLM-based methods fail to effectively decomposed the task, which exacerbates the performance limitations of SLMs in handling complex tasks. Consequently, SLMs have not been effectively applied to NER.
2.3 NER Dataset
@@
A considerable number of NER datasets in various domains have been developed due to the significance of NER Tjong Kim Sang (2002); Kim et al. (2003); Doddington et al. (2004); Walker et al. (2006); Weischedel et al. (2011); Pradhan et al. (2013); Derczynski et al. (2017). However, these existing datasets exhibit several limitations, making them unsuitable for the GEIC task. Most previous multilingual NER datasets adopt coarse-grained classification. For example, PAN-X Pan et al. (2017) includes only three types. Such datasets no longer meet the fine-grained requirements of contemporary flat NER applications. Even existing fine-grained datasets demonstrate clear limitations in category coverage and granularity, falling short of being truly "universal." For instance, FewNERD, despite having 66 entity types, suffers from highly imbalanced data distribution, which affects its reliability for evaluating few-shot learning capabilities. Furthermore, current datasets fail to adequately address the generalization and few-shot/zero-shot learning capabilities of LLMs, hindering the comprehensive training and evaluation of LLM-based NER methods.
3 GEIC
3.1 Definition
Typical NER methods regards NER as a sequence labeling task, where each word is assigned a label, usually BIO representation, and use models to predict entity labels by maximizing the conditional probability of the label sequence:
$P(Y|X)=\prod_{t=1}^{T}P(y_{t}|X,y_{1:t-1})$ (1)
where $X=\{x_{1},\dots,x_{T}\}$ is the input sequence, and $Y=\{y_{1},\dots,y_{T}\}$ is the corresponding label sequence. Such methods focus on local dependencies, making them less effective at capturing sentence-level dependencies. Their reliance on annotated data also limits their adaptability, especially in low-resource scenarios.
@@ 3.1 Definition
Figure 2: Use the sentence "Nothstein beat Buran from Czech Republic" and the multi-granularity categories of AnythingNER as the example. The extractor and classifier are the two different SLMs used in CascadeNER. Azure boxes represent the specific type list for the input of the classifier. Blue boxes represent the sentence input.
@@
Generation-based Extraction and In-context Classification (GEIC), unlike traditional NER tasks, divides NER into two sequential, independently executed, generation-based sub-tasks. In the first sub-task, extraction, the model generate a sentence where all named entities are marked with identifiers and individually re-embeds each entity back into its context, resulting in sentences with identifiers at the number of entities. In the second sub-task, classification, the model receives sentences with identifiers and a list of entity types, and label one entity at a time. When handling multi-granularity data, GEIC allows a hierarchical approach, firstly performing coarse-grained classification and then fine-grained classification based on the identified coarse-grained type and its subcategories.
The unique design of GEIC provides several key advantages. First, GEIC decomposes the NER task into two simpler and independently executed sub-tasks, reducing the complexity of each while enabling the model to better handle diverse and intricate language patterns. This separation allows the sub-tasks to be managed by different models, offering greater flexibility for aggressive optimization strategies that might hinder comprehensive performance. Second, both sub-tasks in GEIC are generation-based and thus well-suited to LLMs. This allows GEIC to leverage the prior knowledge, generalization, and reasoning capabilities of LLMs better, leading to more accurate results, particularly in low-resource scenarios. Third, by performing extraction and classification based on the entire sentence, GEIC aligns with LLMs’ global attention mechanisms, enhancing the relization of contextual relationships. Ultimately, GEIC supports progressive multi-granularity classification, significantly improving performance in handling nuanced and accurate entity classification, making it ideal for the fine-grained requirements of NER applications.
3.2 CascadeNER
@@
To implement GEIC tasks, we propose CascadeNER, a novel framework based on LLMs. A simplified pipeline of CascadeNER with an example is illustrated in Figure 2. To optimize performance while reducing computational resources, CascadeNER employs model cascading, where the extraction and classification sub-tasks are handled separately by two specialized, fine-tuned LLMs, typically SLMs. This structure allows each model to focus on its specific sub-task, maximizing performance on simpler, more specialized tasks and fully aligning with the concept of GEIC.
@@
Upon receiving the input sentence, CascadeNER first processes the sentence by the extractor to mark all named entities with identifiers, and re-embeds each entity back, resulting in sentences with identifiers around the named entities. These sentences are then individually fed into the classifier, which classifies each entity based on the context and the input category list. For multi-granularity categories, CascadeNER employs progressive classification, moving from coarse to fine categories, enhancing its accuracy in handling complex classifications. The detailed steps of extraction and classification are discussed in the following sections.
3.2.1 Extraction

### v4

# DynamicNER: A Dynamic, Multilingual, and Fine-Grained Dataset for LLM-based Named Entity Recognition

@@ Abstract
With the advancement of Large Language Models (LLMs), more and more researchers apply LLMs for Named Entity Recognition (NER) methods, bringing vitality to this classical Natural Language Processing task. However, existing datasets are designed for traditional machine learning methods, inadequate for LLM-based methods in terms of corpus selection, entity categorization, and design logic. This limitation leads to less effective evaluation and model fine-tuning. To address this issue, we propose DynamicNER, the first NER dataset specifically designed for LLMs and with dynamic categorization, transcending the limitations of fixed categorization in existing datasets. It is also multi-lingual and multi-granular, covering 8 languages and 155 entity types, with corpus spanning multiple specialized domains. Furthermore, in response to the limitations demonstrated by existing LLM-based methods during DynamicNER testing, we develop CascadeNER, a novel NER method based on a two-stage strategy and lightweight LLMs, addressing the problems in current methods. Experiments show that DynamicNER is an effective benchmark for LLM-based NER methods, and CascadeNER outperforms existing methods with fewer computational resources. Our work is openly accessible.
DynamicNER: A Dynamic, Multilingual, and Fine-Grained Dataset for LLM-based Named Entity Recognition
Hanjun Luothanks: Zhejiang University, hanjun.21@intl.zju.edu.cn 11footnotemark: 1, Yingbin Jinthanks: Zhejiang University, 3210104412@zju.edu.cn 22footnotemark: 2, Xinfeng Lithanks: Nanyang Technological University, xinfengli@zju.edu.cn 33footnotemark: 3, Xuecheng Liuthanks: Zhejiang University, xuecheng.21@intl.zju.edu.cn 44footnotemark: 4, Ruizhe Chenthanks: Zhejiang University, ruizhec.21@intl.zju.edu.cn 55footnotemark: 5, Tong Shangthanks: Southeast University, 213211756@seu.edu.cn 66footnotemark: 6, Kun Wangthanks: Nanyang Technological University, wk520529@mail.ustc.edu.cn 77footnotemark: 7, Qingsong Wenthanks: Squirrel AI, qingsongedu@gmail.com 88footnotemark: 8, Zuozhu Liuthanks: Zhejiang University, zuozhuliu@intl.zju.edu.cn, Corresponding author 99footnotemark: 9
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) brings revolutions to natural language processing (NLP) (Naveed et al., 2023). As a classical NLP task, Named Entity Recognition (NER) also witnesses significant method innovations driven by LLMs Xie et al. (2023). Existing LLM-based methods Shao et al. (2023); Li and Zhang (2023), leveraging LLMs’ generalization and contextual understanding capabilities, show superior performance compared to traditional machine learning (ML) methods in low-resource, multilingual, or few/zero-shot learning scenarios. This advantage is particularly crucial in AI for Science and Healthcare, where domain resources, especially non-English materials, are often limited. Thus, LLM-based methods gain attention from researchers in these fields Xiao et al. (2024).
Nevertheless, there are no existing NER datasets optimized for the characteristics of LLMs, limiting the effective training and evaluation of LLM-based methods. This limitation manifests primarily in three aspects. First, existing datasets employ static categorization with a fixed set of entity types, preventing effective evaluation of LLMs’ generalization capabilities in handling new entity types and varying granularities under few-shot/zero-shot scenarios. Secondly, the corpus in existing NER datasets focuses on short texts and isolated sentences, making them ineffective to evaluate the capabilities of LLMs in long-range context. Finally, while existing datasets separately address specialized corpus with corresponding entity types Kim et al. (2003); Liu et al. (2021), multi-grained classifications Ding et al. (2021), and multi-languages Malmasi et al. (2022), no dataset simultaneously incorporates three aspects, limiting comprehensive evaluation of LLM-based methods, which show potential across all these aspects. As a result, existing datasets lack the capability to effectively discriminate performance differences among LLM-based methods, fail to reflect their true potential and limitations, and impedes future improvements.
To address this issue, we develop DynamicNER, the first NER dataset optimized for LLM-based methods and the first NER dataset with dynamic categorization. It employs multiple strategies to dynamically adjust entity labels, type lists, and granularity levels during annotation, enhancing the evaluation of generalization ability of NER methods. We design cohesion and distribution balance metrics to enable evaluation and optimization during the categorization, with the entire procedure being algorithmically automated, ensuring reliability and reproducibility. This method addresses the limitations of existing datasets in training and evaluating few-shot/zero-shot learning ability. Besides, DynamicNER is a multilingual and multi-granular dataset, featuring 8 languages, 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types, with entity types and corpora covering extensive professional content across domains including sciences, medicine, and art, providing an unprecedented level of coverage.
Furthermore, our test on DynamicNER reveals significant limitations in existing methods, particularly when migrating to lightweight LLMs (LLMs with parameters from 1.5B to 7B) for local deployment. While methods with commercial models like GPT Brown (2020) achieve reported performance levels, this dependency introduces challenges regarding API costs and privacy risks. Using models via API are expensive for practical NER application. Privacy concern is also a significant issue Zhang et al. (2024); Das et al. (2024); Deng et al. (2025). Given the current absence of laws regarding LLM data uploading in many countries, users face difficulties in protecting their rights. Therefore, preventive measures to avoid unauthorized disclosure of information becomes essential.
For this issue, we propose CascadeNER, a universal and multilingual NER framework that achieves comparable performance with lightweight LLMs to existing LLM-based methods requiring expensive commercial models. CascadeNER employs a two-stage strategy, redefines NER as two in-context text generation sub-tasks, extraction and classification. By model cascading Varshney and Baral (2022), CascadeNER assigns each sub-tasks to corresponding fine-tuned lightweight LLMs, decreasing task complexity and helping capture in-context dependencies. Leveraging the prior knowledge, CascadeNER can complete NER tasks in multiple languages in low-resource scenarios effectively.
We evaluate a BERT-based Devlin et al. (2018) supervised method, two LLM-based methods, and our proposed CascadeNER on DynamicNER. The three LLM-based methods are tested on 3 difference models: Qwen2.5-1.5B, Qwen2.5-7B Bai et al. (2023), and GPT-4o Achiam et al. (2023). We also conduct evaluations of CascadeNER against existing methods on existing datasets. Results demonstrate that DynamicNER effectively evaluates the performance of LLM-based methods in low-resource and complex NER tasks, while CascadeNER outperforms existing LLM-based methods significantly with smaller models. Moreover, this work offers the first comprehensive comparison and analysis of existing LLM-based NER methods, with a emphasis on multilingual and fine-grained scenarios.
@@
  • We build DynamicNER, the first NER dataset optimized for LLM-based NER with a novel dynamic categorization system, which supports 8 languages, 155 entity types, and three-level granularity
  • We propose CascadeNER, a universal and multilingual NER framework, which outperforms existing LLM-based methods with lightweight LLMs
  • We conduct the first comprehensive comparison of LLM-based NER methods, discussing future directions of LLM-based NER
@@ 2 Related Works
Named Entity Recognition
@@
NER is the task of identifying named entities in text and classifying them into predefined categories. Supervised methods, such as BiLSTM Yu et al. (2020) and BERT-MRC Li et al. (2019a), currently dominate this task. They generally rely on large amounts of train data to achieve strong performance, which limits their application in low-resource scenarios. Some researchers apply LLMs to address this issue. GPT-NER Wang et al. (2023) employs the GPT-3 model and re-frames the task as single-entity labeling, supporting few-shot/zero-shot learning. It achieves comparable performance to supervised methods in traditional scenarios and excels in low-resource scenarios. PromptNER Ashok and Lipton (2023) achieves state-of-the-art (SOTA) accuracy in datasets with complex classification Liu et al. (2021); Ding et al. (2021) with GPT-4 and Chain-of-Thought (CoT) Wei et al. (2022), yet performs significantly worse than GPT-NER and supervised methods in classical NER datasets like CoNLL2003 Tjong Kim Sang and De Meulder (2003). Furthermore, several studies apply LLM-based NER in domain-specific tasks Li and Zhang (2023); Shao et al. (2023); Keloth et al. (2024), focusing on science and medicine. Their performances surpass supervised methods in those domains, further highlighting the potential of LLMs in low-resource and complex NER tasks.
NER Dataset
@@
There have been a considerable number of NER datasets in various domains Tjong Kim Sang (2002); Kim et al. (2003); Doddington et al. (2004); Walker et al. (2006); Weischedel et al. (2011); Pradhan et al. (2013); Derczynski et al. (2017). However, these existing datasets exhibit several limitations, making them unsuitable for LLM-based NER. Most previous multilingual NER datasets adopt coarse-grained classification, no longer meeting the fine-grained requirements of contemporary flat NER applications. Even existing fine-grained datasets demonstrate clear limitations in category coverage and granularity, falling short of being truly "universal." For instance, FewNERD, despite having 66 entity types, suffers from highly imbalanced data distribution, which affects its reliability for evaluating few-shot learning capabilities. Furthermore, current datasets fail to adequately address the generalization capabilities of LLMs, hindering the comprehensive training and evaluation of LLM-based NER methods. Table 1 presents a simple comparison between DynamicNER and existing multilingual or fine-grained datasets.
Dataset #Language #Coarse #Fine Domain
CoNLL2002 2 4 no News
CoNLL2003 2 4 no News
ACE2005 2 7 41 News
OntoNotes 5.0 3 18 no General
CrossNER 1 9-17 no Multi Domain
FEW-NERD 1 8 66 General
PAN-X 282 3 no General
MultiCoNER 11 6 33 General
I2B2 1 22 no Medical
DynamicNER 8 8 155 Multi Domain
Table 1: Overview of NER datasets. Notably, DynamicNER covers a wide range of cross-domain categories, such as art, medicine, and biology, thus offering better generalization compared to other general datasets.
3 DynamicNER Dataset
DynamicNER spans 8 languages: English, Chinese, Spanish, French, German, Japanese, Korean, and Russian. In terms of categorization, it is the first NER dataset with three-level granularity categorization, encompassing 8 coarse-grained types, 31 medium-grained types, and 155 fine-grained types, as shown in Figure 1. Like other NER datasets, DynamicNER is divided into train, dev, and test sets, and data volumes for different languages and parts shown in Appendix C. To develop DynamicNER, we first collect unlabeled corpus from Wikipedia. Then we manually extract sentences from corpora and annotate entities. After human annotation, we guide the dynamic categorization with category cohesion and distribution uniformity to automatically process base DynamicNER, and results in one base version and one dynamic version. Details are given in the following parts.
Figure 1: The coarse-grained and medium-grained categories of DynamicNER. Detailed categories are provided in Appendix J.
Corpora Collection and Annotation
Wikipedia provides multilingual, domain-specific corpora with clear hierarchical and indexing systems, serving as a rich resource for our research. We utilize legal Wikipedia-API to filter and download corpora across different languages and categories, followed by manual selection and annotation of sentences. We particularly focus on corpora containing long texts and complex contexts. After completing 50% of the annotation process of each language, we annotate corpora from categories related to underrepresented entity types to achieve a balanced entity type distribution. For instance, when the entities of "Algorithm" are significantly less than others, we use more corpora from Computer Science category. As a result, DynamicNER ensures balanced entity distribution, and includes rare entities and emerging fields which are not adequate in existing datasets, ensuring comprehensive coverage across diverse domains. A statement about annotators and ethical considerations is provided in Appendix I.
Dynamic Categorization
The dynamic version improves model generalization and reduces overfitting risk by dynamically adjusting entity labels and corresponding entity type lists during annotation, including mixing types of different granularities, replacing types with synonyms, using type lists without irrelevant types, and merging certain types into miscellaneous/others, as shown in Figure 2. This method addresses the mismatch between existing datasets and few-shot/zero-shot training needs, better simulating real-world scenarios, and is particularly critical for evaluating methods relying on complex prompt designs (e.g., CoT). Unlike traditional few-shot learning, some LLM-based methods only use few-shot demos to help the model understand the task or format, without requiring knowledge of entity types. They can perform NER across different datasets with fixed few-shot demos, resembling zero-shot NER. Research shows this method is more effective than typical zero-shot NER Zhang et al. (2022). In methods that uses complex prompt designs like CoT to guide the reasoning, even few-shot CoT only conveys the CoT process rather than task-relevant knowledge, the performance of prompt-guide zero-shot CoT is significantly worse than few-shot CoT, making zero-shot restrictions inadequate for reflecting their true capability. However, in NER, models inevitably learn about entity types through few-shot demos, which limits generalization evaluations on fixed-category datasets. Our method significantly mitigates this limitation by varying entity types and lists, isolating the impact of prior type knowledge.
Figure 2: Examples of dynamic categorization. The red box contains the fine-grained categories in base version. Green boxes contains the lists modified with 4 methods. Blue boxes contains the entity type to each list.
Notably, as dynamic categorization is a subtractive method applied to a comprehensive classification system, this method highly relies on DynamicNER’s comprehensive categorization system, which includes 155 entity types. This method may not be suitable for all datasets.
Categorization Metrics
Random dynamic categorization not only exhibits poor reproducibility and explainability, but may also lead to data quality degradation. For training, inappropriate categorization may result in inconsistent learning objectives and overfitting risks Ren et al. (2016). For evaluation, certain categories may experience imbalanced sample distribution and boundary ambiguity, reducing the comprehensiveness and consistency of evaluation Obeidat et al. (2019). Thus, we design four metrics to regulate the dynamic categorization: cohesion, normalized entropy, Gini coefficient, and variation coefficient. The definition and calculation methods are provided in Appendix A.
Figure 3: Pipeline of dynamic categorization. Blue boxes contains methods of each re-categorization. Yellow boxes contains used metrics of each evaluation.
Categorization Process
The dynamic categorization process consists of 4 rounds of re-categorization, each sequentially corresponding to an adjustment method, and different metrics are employed in each round to guide the optimization. This hierarchical design enables each stage to focus on distinct data characteristics and optimization objectives, preventing interference between metrics while ensuring proper optimization direction, thus achieving a progressive optimization. We do not use all metrics in each evaluation, considering that certain metrics may have overlapping or conflicting effects at specific stages. For instance, normalized entropy and Gini coefficient both measure distribution uniformity, while improving cohesion may lead to more concentrated distribution and consequently lower entropy values. Figure 3 illustrates the metrics and methods corresponding to each round. Appendix B explains the reasons of metric selection for each round.
4 CascadeNER
@@
Figure 4: Use a sentence and the multi-granularity categories of DynamicNER as the example. The extractor and classifier are the two different lightweight LLMs used in CascadeNER. Azure boxes represent the specific type list for the input of the classifier. Blue boxes represent the sentence input.
4.1 Framework
Background
Some existing supervised methods suggest that separating extraction and classification can improve NER performance as this two-stage strategy reduces the task complexity Shen et al. (2021); Wu et al. (2022). However, these methods are limited by traditional models, failing to incorporate LLMs, and exhibit notable performance deficiencies that make them inferior to other methods treating NER as a single task. On the other hand, LLM-based methods demonstrate superior performance compared to traditional methods in Named Entity Extraction Sancheti et al. (2024) and Text Classification Gasparetto et al. (2022), indicating the potential of two-stage in LLM-based NER.
Framework Design
@@ 4.1 Framework
We propose the framework to implement two-stage strategy in LLM-based NER. CascadeNER divides NER into two sequential, independently executed, generation-based sub-tasks. In the first sub-task, extraction, the model generate a sentence where all named entities are marked with identifiers and individually re-embeds each entity back into its context, resulting in sentences with identifiers at the number of entities. In the second sub-task, classification, the model receives sentences with identifiers and a list of entity types, and label one entity at a time.
Model Cascading
@@
To optimize performance while reducing computational resources, CascadeNER employs model cascading, where the extraction and classification sub-tasks are handled separately by two specialized fine-tuned LLMs. This structure allows each model to focus on its specific sub-task, maximizing performance on simpler, more specialized tasks. The architecture enables CascadeNER to be particularly suitable for lightweight LLMs, as each model only focus on a simplified task. Existing research shows that fine-tuned lightweight LLMs can achieve performance close to normal LLMs on specific simple tasks Hu et al. (2024a). Through the implementation of two-stage strategy and model cascading, CascadeNER effectively leverages the advantages of lightweight LLMs in simple tasks, maintaining high accuracy while reducing computational resource usage.
Pipeline
@@
A simplified pipeline of CascadeNER with an example is shown in Figure 4. Upon receiving the input sentence, CascadeNER first processes the sentence by the extractor to mark all entities with identifiers, and re-embeds each entity back, resulting in sentences with identifiers around the named entities. These sentences are then individually fed into the classifier, which classifies each entity based on the context and the input type list. For multi-granularity data, CascadeNER allows a progressive strategy, significantly improving CasacdeNER’s performance in accurate fine-grained classification. The detailed steps of extraction and classification are discussed in following sections.
4.2 Extraction
Prompt Design